OpenStudy (anonymous):

-x+2y=2 write the equation using functional notation

OpenStudy (anonymous):

Solve for y and replace it with f(x). -x+2y=2 2y=x+2 y=(1/2)x+1 f(x)=(1/2)x+1

OpenStudy (anonymous):

thanks, I got that right but don't understand how you got f()=(1/2)x+1